How many hydrogen atoms are in
8.50 mol
8.50 mol
of ammonium sulfide?
1
Expert's answer
2019-07-19T07:53:43-0400

First, the number of ammonium sulfide molecules should be calculated:

N = NA × n,

where NA - the Avogadro number, n - number of moles.

N (ammonium sulfide) = 6.022 × 1023 × 8.5 mol = 51.187 × 1023.

The moelcular formula of ammonium sulfide is (NH4)2S. It means that each molecule contains 8 hydrogen atoms.

As a result, 8.5 mol of (NH4)2S contain:

51.187 × 1023 × 8 = 41 × 1024 hydrogen atoms.

Answer: 41 × 1024 hydrogen atoms